OpenStudy (anonymous):

On discovering that her family had a 70% risk of heart attack, Erin took a treadmill test to check her own potential of having a heart attack. The doctors told her that the reliability of the stress test is 67%. What is the probability that Erin will have a heart attack and the test predicts it? A. 0.099 B. 0.201 C. 0.231 D. 0.469

OpenStudy (kirbykirby):

I'm not 100% sure, but say \(H\) = potential for heart attack \(T\) = test predicts it Then you are given \( P(H)=0.7\) \(P(T|H) = 0.67\), I believe that reliability is saying what the probability of test predicting it given that you have a potential for heart attack So then you are asked to find \(P(T \cap H)\). \[P(T|H)=\frac{P(T \cap H)}{P(H)} \implies P(T \cap H)=P(T|H)P(H)=0.67*0.7\]